### v2

# ESGenius: Benchmarking LLMs on Environmental, Social, and Governance (ESG) and Sustainability Knowledge

Abstract
We introduce ESGenius, a comprehensive benchmark for evaluating and enhancing the proficiency of Large Language Models (LLMs) in Environmental, Social, and Governance (ESG) and sustainability-focused question answering. ESGenius comprises two key components: (i) ESGenius-QA, a collection of 1,136 Multiple-Choice Questions (MCQs) generated by LLMs and rigorously validated by domain experts, covering a broad range of ESG pillars and sustainability topics. Each question is systematically linked to its corresponding source text, enabling transparent evaluation and supporting Retrieval-Augmented Generation (RAG) methods; and (ii) ESGenius-Corpus, a meticulously curated repository of 231 foundational frameworks, standards, reports, and recommendation documents from 7 authoritative sources. Moreover, to fully assess the capabilities and adaptation potential of LLMs, we implement a rigorous two-stage evaluation protocol—Zero-Shot and RAG. Extensive experiments across 50 LLMs (0.5B to 671B) demonstrate that state-of-the-art models achieve only moderate performance in zero-shot settings, with accuracies around 55–70%, highlighting a significant knowledge gap for LLMs in this specialized, interdisciplinary domain. However, models employing RAG demonstrate significant performance improvements, particularly for smaller models. For example, DeepSeek-R1-Distill-Qwen-14B improves from 63.82% (zero-shot) to 80.46% with RAG. These results demonstrate the necessity of grounding responses in authoritative sources for enhanced ESG understanding. To the best of our knowledge, ESGenius is the first comprehensive QA benchmark designed to rigorously evaluate LLMs on ESG and sustainability knowledge, providing a critical tool to advance trustworthy AI in this vital domain.
ESGenius: Benchmarking LLMs on Environmental, Social, and Governance (ESG) and Sustainability Knowledge
Chaoyue He1   Xin Zhou1 thanks: Corresponding author  Yi Wu1   Xinjia Yu1   Yan Zhang1   Lei Zhang1   Di Wang1 Shengfei Lyu1Hong Xu1Xiaoqiao Wang2Wei Liu2Chunyan Miao1 1Alibaba-NTU Global e-Sustainability CorpLab (ANGEL), Singapore;   2Alibaba Group, China {cyhe, xin.zhou, xinjia.yu, zhang.yan, wangdi, shengfei.lyu, xuhong, ascymiao}@ntu.edu.sg {wuyi0614, leizhanzzl.1103}@gmail.com   {nerissa.wxq, weiliu.liuwei}@alibaba-inc.com GitHub: https://github.com/ANGEL-NTU/ESGenius Web Portal: https://angel-ntu.github.io/ESGenius
1 Introduction
Figure 2: Relationship between model size and zero-shot accuracy across 50 LLMs evaluated on the ESGenius benchmark. Model sizes are plotted on a $\log_{10}$ scale (in billions of parameters), with accuracy shown as percentages. There is a moderate positive correlation between model size and performance, suggesting larger models generally perform better. Dot colors denote five model families (DeepSeek, Gemma, GPT, Meta-Llama, Qwen). For proprietary API models, parameter counts are based on industry estimates (See Table 2 for details).
Environmental, Social, and Governance (ESG) knowledge encompasses a vast domain of sustainability and corporate responsibility information that Large Language Models (LLMs) must effectively process to serve emerging business needs Zhou et al. (2024); Singh et al. (2025); Ong et al. (2025); Zou et al. (2025); Zhang et al. (2024); Yu et al. (2025). This field spans critical areas from climate change and emissions tracking IPCC (2023) to workplace safety and human rights Global Reporting Initiative (2023). Such knowledge is codified in technical frameworks and standards—major ones including GRI Initiative (2023), SASB Board (2023), TCFD Board (2017), ISSB International Sustainability Standards Board (2023), and CDP CDP Worldwide (2023)—all of which are constantly evolving.
While LLMs show promise in processing complex ESG documents and providing relevant answers to user queries, their capabilities in this interdisciplinary domain remain largely unevaluated. Considering the high stakes involved, where incorrect responses about ESG requirements or metrics could lead to serious compliance violations or misguided sustainability initiatives, this assessment gap is particularly problematic. However, there is currently no comprehensive question-answering (QA) benchmark specifically designed to evaluate how well LLMs understand and reason about ESG concepts. Existing QA benchmarks either omit ESG topics entirely or address them only superficially. This gap leaves researchers and practitioners without a reliable way to measure and improve LLMs’ ESG knowledge comprehension and question-answering abilities.
To close this gap, we present ESGenius, a curated benchmark that targets Multiple-Choice Questions (MCQs) answering as the core evaluation task for ESG and sustainability knowledge (Figure LABEL:fig:esgenius_pipeline). Our contributions are fourfold: (1) ESGenius benchmark: A comprehensive evaluation framework comprising two integrated components: (i) ESGenius-QA, a collection of 1,136 MCQs across various ESG pillars and sustainability topics, generated using state-of-the-art (SOTA) LLM approaches and validated by domain experts. Each question is explicitly mapped to supporting evidence from authoritative source texts, enabling transparent evaluation and facilitating RAG applications; and (ii) ESGenius-Corpus, a collection of 231 ESG documents and frameworks enabling efficient knowledge retrieval from 7 major authoritative sources. (2) Evaluation Protocol: We implement a comprehensive two-stage evaluation protocol consisting of zero-shot testing and RAG to systematically assess LLM capabilities. This assessment framework provides valuable insights into the current limitations and future potential of LLMs in understanding ESG. (3) Evaluation Analysis: Testing across a diverse set of 50 LLMs (0.5B to 671B) reveals significant performance gaps in zero-shot settings, with most models achieving 55–70% accuracy and the best model (o3) achieving a top score of 72.54%, as shown in Figure 2. However, models demonstrate substantial potential for improvement through RAG approaches, with DeepSeek-R1-Distill-Qwen-14B improving from 63.82% (zero-shot) to 80.46%. (4) Open Source Initiative: To foster community engagement and collaborative advancement, we have made our complete benchmark suite publicly available at https://github.com/ANGEL-NTU/ESGenius. This includes comprehensive documentation, evaluation code, model implementations, and the full ESGenius dataset. We also maintain an interactive web portal at https://angel-ntu.github.io/ESGenius featuring a real-time leaderboard and detailed performance visualizations through heatmap and one of its cells (Figures 9 and 10), enabling researchers to track progress and identify areas for improvement in ESG-focused language models.
The remainder of this paper is organized as follows: §2 surveys prior efforts on ESG knowledge resources, question-answering benchmarks, and retrieval-augmented generation; §3 details the construction of the ESGenius benchmark, describing both the ESGenius-QA dataset and the ESGenius-Corpus; §4 outlines our experimental protocols—including zero-shot and RAG settings—and presents a comprehensive evaluation of 50 LLMs; §5 concludes the paper and highlights directions for subsequent research; §6 discusses the limitations of our benchmark, while §7 reflects on ethical considerations.
2 Related Work
Prior work at the intersection of LLMs and ESG has spanned diverse strands of research, from building domain-specific corpora and ontologies to designing evaluation benchmarks and retrieval-augmented generation methods. In this section, we review relevant literature across three key areas: (i) ESG and sustainability knowledge resources, which provide the taxonomies, corpora, and specialized models needed to capture domain-specific terminology; (ii) QA and evaluation benchmarks, which illustrate how standardized datasets have driven progress in measuring factual knowledge, reasoning, and domain expertise, while also exposing the absence of ESG-focused QA benchmarks; and (iii) RAG and knowledge grounding, which highlight how retrieval mechanisms enhance interpretability and factual accuracy in specialized domains such as finance and climate science.
2.1 ESG & Sustainability Knowledge Resources
ESG data have traditionally been guided by voluntary reporting frameworks such as the Global Reporting Initiative (GRI) and the Sustainability Accounting Standards Board (SASB). These frameworks define taxonomies of ESG topics (e.g., emissions, labor practices, board governance) and indicators, but the unstructured text in corporate sustainability reports poses significant challenges for computational use. To impose structure, researchers have proposed ontologies and knowledge bases for ESG. For instance, Zhou and Perzylo (2023) introduced OntoSustain, aligning GRI and EU standards (ESRS) to capture key sustainability indicators, though broad coverage and automated population remain difficult.
Several ESG-focused text corpora have emerged to support NLP research in this domain. One recent example is SusGen-3K, a dataset of 30k instances spanning multiple financial NLP tasks (e.g., sentiment, Q&A) and an ESG report generation task Wu et al. (2025). Meanwhile, Chang et al. (2024) built a benchmark of more than 8,000 labeled sentences from more than 14,000 corporate sustainability reports covering 36 ESG topics for topic and quality classification. Other datasets focus on more specific areas, such as the benchmark from Beck et al. (2025) for extracting greenhouse gas emissions, or on non-English languages, like the ESG-Kor dataset for information extraction in Korean Lee et al. (2024). These datasets remain relatively restricted, especially compared to general-domain corpora, and focus primarily on classification or generation tasks rather than MCQs question answering, thus motivating the need for broader coverage like ESGenius.
Prior NLP work on sustainability texts has largely centered around information extraction and text classification, evolving from early methods that relied on topic models or lexicon-based approaches Raghupathi et al. (2020) but often fell short in capturing nuanced ESG terminology. More recent efforts leverage LLMs, either by developing specialized models or by applying general models in sophisticated pipelines. In the former category, researchers have introduced domain-specific models like ClimateBERT, which continued BERT’s pre-training on 1.6M climate-related paragraphs for more effective risk disclosure classification Webersinke et al. (2021), and E-BERT, a set of models designed to optimize ESG reporting Zhang et al. (2025eb). In the latter category, approaches like ChatReport use GPT-based summarization and semantic search to assess a company’s report against TCFD recommendations Ni et al. (2023), while others have employed GPT-based pipelines to extract triples (company, action, ESG issue) and populate knowledge graphs from sustainability disclosures Bronzini et al. (2024). Collectively, these advanced approaches illustrate the promise of LLMs for ESG tasks but also highlight the need for standardized benchmarks for a thorough evaluation.
2.2 QA & Evaluation Benchmarks
Question answering (QA) benchmarks have driven many NLP advancements over the past decade, starting with the TREC QA tracks Voorhees et al. (1999) and continuing through datasets like SQuAD Rajpurkar et al. (2016). While these resources predominantly test general factual or reading comprehension, specialized QA datasets have emerged to measure higher-level reasoning and domain expertise. For example, OpenBookQA Mihaylov et al. (2018) challenged models to combine a small “open book” of scientific facts with common sense, while BoolQ Clark et al. (2019) featured yes/no queries requiring nuanced text understanding. Datasets such as DROP Dua et al. (2019) emphasize discrete reasoning skills, involving counting or arithmetic over paragraphs.
Further expansions have produced large-scale evaluation suites such as MMLU (Massive Multitask Language Understanding) Hendrycks et al. (2020) and MMMU (Massive Multi-Discipline Multimodal Understanding) Yue et al. (2024), which cover dozens of subjects at the college-exam level, along with domain-specific benchmarks like PubMedQA Jin et al. (2019) for biomedical research, JEC-QA Zhong et al. (2020) for legal question answering, and FinanceBench Islam et al. (2023) for finance. Yet despite these advances, there is still a notable lack of QA benchmarks dedicated to ESG and sustainability. To address this gap, ESGenius introduces a large collection of expert-validated MCQs spanning environmental, social, and governance topics. Drawing from authoritative frameworks (e.g., IPCC, GRI, SASB), ESGenius provides a high-stakes, real-world testbed tailored to ESG.
2.3 RAG and Knowledge Grounding
When tackling specialized or dynamically updated domains, RAG has emerged as a powerful method for accurate and interpretable QA. Lewis et al. (2020) introduced RAG, an approach that combines a parametric language model with a non-parametric memory of documents. This approach improved performance on knowledge-intensive QA tasks relative to purely parametric models, as it allows the LLM to ground responses in retrieved evidence.
Other retrieval-generation frameworks such as GraphRAG Edge et al. (2024), Fusion-in-Decoder Izacard and Grave (2021) and REALM Guu et al. (2020) illustrated diverse strategies for fetching relevant text chunks to guide the model’s reasoning. Recently, these methods have been applied to high-stakes sectors including finance and climate science. For instance, Vaghefi et al. (2023) integrated GPT-4 with a 3,000-page IPCC report to answer climate change questions more accurately. Similarly, Ni et al. (2023) used semantic search to retrieve TCFD-related segments from a company’s sustainability report, prompting an LLM to summarize potential climate risks. This grounding is critical in ESG contexts, where traceability and authoritative references are essential.
3 The ESGenius Benchmark
Developing reliable AI tools for the complex ESG and sustainability domain requires high-quality, domain-specific data. Our ESGenius benchmark provides a comprehensive foundation to evaluate and enhance LLMs in this context, drawing from globally recognized standards and authoritative sources. As illustrated in Figure LABEL:fig:esgenius_pipeline, the benchmark is built through a structured pipeline where ESG-related documents are digitized, LLMs generate candidate MCQs, and domain experts validate them to form ESGenius-QA, while the underlying source materials are systematically curated into the ESGenius-Corpus. Together, these two interconnected components ensure explicit linkages between questions and textual evidence, enabling transparent evaluation of ESG knowledge and supporting RAG methods. The detailed coverage of sources and statistics is presented in Appendix D.
3.1 ESGenius-QA
QA Generation and Validation Principles.
The ESGenius-QA dataset comprises 1,136 carefully curated MCQs with ground truth answers, validated by ESG domain experts. The questions systematically cover essential ESG topics across three core pillars including but not limited to: (i) Environmental: Climate change mitigation and adaptation, comprehensive carbon accounting (Scope 1, 2, 3 emissions), energy efficiency, water resource management, biodiversity conservation, waste reduction, and pollution control, etc; (ii) Social: Labor practices and standards, human rights protection, diversity and inclusion initiatives, workplace health and safety, community engagement and impact assessment, data privacy and protection, etc; (iii) Governance: Board structure and independence, executive compensation frameworks, business ethics and compliance, shareholder rights and engagement, enterprise risk management, and regulatory adherence, etc. The dataset is generated through our specialized pipeline (Appendix A.1) and undergoes rigorous validation by domain experts to ensure accuracy and relevance (§3.1).
Descriptions.
The ESGenius-QA dataset is structured to reflect the complexity and diversity of real-world ESG assessments through a standardized multiple-choice format. Each question contains four options (A-D) with a single correct answer, plus a dedicated “Not sure” option (Z) to capture model uncertainty. This design enables systematic evaluation of both performance and confidence. The dataset carefully balances conciseness with domain-specific precision. Questions and options employ precise ESG terminology while maintaining clarity and focus. Contextual prompts are provided where necessary for essential background. The specialized lexicon throughout authentically represents real-world ESG assessment scenarios. Table 1 presents a detailed token-level analysis of questions, options, and source text. Key terminology distributions are visualized through word clouds for questions (Figure 4), answer options (Figure 5) and the source text (Figure 6).
Corpus-Question Mapping.
Each question in ESGenius-QA is precisely mapped to relevant page(s) and text passages within the 7 authoritative source documents (GRI, SASB, IPCC, etc.) containing the information required for correct answers. ESG domain experts meticulously validate these mappings, ensuring both accuracy and relevance. This explicit linking between questions and their supporting evidence in source texts enables effective RAG approaches.
QA Quality Control.
To ensure the highest quality of questions and answers, we employ a multi-stage validation process (Appendix A.1). While Qwen-max-2025-01-25 was used in the initial phase to generate candidate QA pairs, the final benchmark was entirely curated by human experts: six independent reviewers with an average of five years’ experience in sustainability or NLP double-validated every question-answer pair, supported by three consulting industry experts with over ten years of relevant experience. Validators confirmed factual correctness against authoritative sources, clarity and unambiguity, plausibility yet incorrectness of distractors, and adequate context. Any question failing these criteria was refined or discarded, yielding a rejection rate of about 25% (1,136 of 1,519 MCQs retained). Importantly, Qwen-max-2025-01-25 was later evaluated only through its official API in a fully isolated session, ensuring no overlap or leakage between construction and evaluation. This expert-in-the-loop curation decisively eliminates circular dependency, and, consistent with standard practices in high-impact benchmarks, our method combines LLM-assisted generation with expert oversight to balance scale, quality, and domain relevance, while remaining open to validated alternatives to further reduce bias.
3.2 ESGenius-Corpus
The ESGenius-Corpus underpins every evaluation in this work, unifying authoritative frameworks, corporate questionnaires, and peer-reviewed scientific assessments that span the full breadth of environmental, social, and governance (ESG) concerns. A quantitative snapshot of the collection appears in Table 3 for IPCC climate-change assessments, Table 4 for the GRI Standards, Tables 5 and 6 for SASB’s standards, Tables 7 and 8 for IFRS/ISSB guidance, Table 9 for TCFD materials, Table 10 for CDP questionnaires, and Table 11 for the UN Sustainable Development Goals (SDGs). These statistics confirm (i) broad topical coverage across all three ESG pillars, (ii) deep sector-level granularity via GRI and SASB industry standards, and (iii) an expanding climate-finance focus through IFRS-ISSB, TCFD, and CDP additions.
Sourcing and Collection.
The corpus integrates sources along three tiers: (1) Core reporting standards include the Global Reporting Initiative (GRI) Standards Initiative (2023), Sustainability Accounting Standards Board (SASB) Standards Board (2023), and International Sustainability Standards Board (IFRS-ISSB) Standards International Sustainability Standards Board (2023); (2) Specialized reports and frameworks extend coverage with IPCC Assessment Reports IPCC (2023), TCFD Guidelines Board (2017), and CDP questionnaires CDP Worldwide (2023); and (3) Overarching global targets are provided by the UN SDGs United Nations (2015).
Extensibility.
The architecture of the ESGenius-Corpus is deliberately modular, enabling the rapid integration of new materials such as emerging regulations (e.g., EU CSRD, ISSB updates), sector-specific guidance (e.g., sustainable finance taxonomies), country-level policies, corporate ESG and sustainability reports, and certification frameworks (e.g., LEED for sustainable buildings). This forward-looking design ensures that the benchmark can evolve in step with the shifting ESG landscape, safeguarding its long-term relevance and maintaining ESGenius as a reliable yardstick for tracking LLM progress in sustainability understanding.
Copyrights and Privacy Handling.
The ESGenius-Corpus comprises publicly available documents and open-access materials. For proprietary standards and frameworks (e.g., GRI, SASB, CDP), we provide references to their official source locations, allowing users to access them directly, with only brief excerpts referenced for academic purposes. The corpus documentation explicitly lists included sources and provides detailed guidance on accessing external materials, ensuring reproducible research while adhering to intellectual property rights. To protect privacy, the corpus excludes all personal and sensitive information through a rigorous manual auditing process, conducted alongside the source mapping step (§3.1) and the quality control stage for dataset curation (§3.1).
4 Experiments and Results
Through ESGenius, we benchmark 50 LLMs spanning open-source checkpoints—DeepSeek Guo et al. (2025), Meta-Llama 3 Grattafiori et al. (2024), Google Gemma 3 Team et al. (2025), and Alibaba Qwen 2.5/3 Yang et al. (2024, 2025) from Hugging Face Hugging Face (2024) and Transformers library Wolf et al. (2020)—as well as proprietary APIs (GPT, DeepSeek, Qwen). All experiments run on a DGX node (4 $\times$ 80 GB A100 GPUs) with fixed random seeds (seed=42) for reproducibility. Our evaluation suite comprises two progressively stronger settings: (1) Zero-Shot Prompting4.1.1), which probes the intrinsic ESG knowledge encoded during generic pre-training, and (2) Long-Context RAG4.1.2), a RAG baseline that prepends relevant evidence to each prompt as long context. The RAG experiments are conducted on 43 open-source models.
For each model-setting combination, we compute exact-match accuracy on single-answer questions. Because LLMs often generate raw outputs that may not directly match the required format, we apply response validation to ensure valid multiple-choice answers (Appendix A.2). The results are presented as a comprehensive leaderboard (Appendix C, Table 2) for both zero-shot and RAG settings. To complement these results, we also include a ranking bar plot in Figure 3, providing a clearer visualization of LLM performance under the zero-shot setup.
4.1 Evaluation Protocols
To rigorously assess model performance on ESGenius, we design two complementary evaluation protocols. The first, Zero-Shot Prompting, measures how well models can answer ESG-related questions without external assistance, reflecting their intrinsic knowledge and reasoning ability. The second, Long-Context RAG, evaluates models in a retrieval-augmented setting by providing each question with its corresponding evidence passage from the ESGenius-Corpus. Together, these protocols enable a transparent comparison between raw model capabilities and knowledge-grounded reasoning, thereby establishing fair and reproducible baselines for future research.
4.1.1 Zero-Shot Prompting
Inference Protocol.
Each model is provided the zero-shot prompt template (Appendix B.1) and an zero-shot question example is shown in Appendix A.3.1. Response generation employs greedy decoding (temperature=0, top_p=1, top_k=0), with a maximum sequence length of 1024 tokens.
Implementation.
Models are evaluated in half precision (FP16) using standard batched inference on GPUs. Results are logged incrementally with error handling to ensure robustness.
4.1.2 Long-Context RAG
Retrieval Protocol.
For each question, the pre-linked source passage from ESGenius-QA is retrieved from the ESG knowledge corpus. This passage is prepended to the prompt template (Appendix B.2), and an illustrative example is provided in Appendix A.3.2. This simple but effective retrieval ensures that models consistently access relevant domain knowledge during inference.
Implementation.
The evaluation pipeline mirrors the zero-shot setup, with the key difference being the augmented prompts. Context is injected by prepending the retrieved passage to each question prompt before feeding it to the model, using the same decoding parameters (temperature=0, top_p=1) and evaluation metrics. Notably, even this straightforward lexical-matching long-context RAG method yielded measurable performance gains, underscoring the benchmark’s utility. While more sophisticated retrieval-augmented generation techniques could further improve performance, our aim was to establish a transparent, fair, and reproducible RAG baseline.
4.2 Main Results
Drawing from the experimental results presented in Table 2, we highlight three key findings and provide detailed analysis from various perspectives subsequently.
4.2.1 Key Findings
(1) ESG Concept Understanding Remains Challenging for Current LLMs. Our evaluation reveals that the top-performing zero-shot model, o3, attained an accuracy of 0.7254, whereas the majority of other models scored approximately 0.65. This performance level is notably comparable to, or lower than, that observed in benchmarks from other domains Guha et al. (2023), underscoring the challenges current LLMs face in accurately comprehending and reasoning about complex ESG concepts. Further details are provided in §4.2.2.
(2) Domain-Specific RAG Allows Smaller Models to Outperform Larger Zero-Shot Counterparts. Performance evaluation demonstrates that applying RAG significantly enhances the capabilities of smaller models on ESG question answering. For example, the accuracy of Qwen3-4B increased from 0.6188 (zero-shot) to 0.7905 (RAG). Similarly, Qwen2.5-1.5B rose from 0.5484 (zero-shot) to 0.6972 (RAG). These improvements demonstrate that targeted retrieval is more effective than raw model scale for complex domain-specific knowledge. More details are described in §4.2.3.
(3) Reasoning Capabilities Enhance LLM Performance on ESG Understanding. Models augmented with explicit reasoning capabilities demonstrate enhanced performance on ESG tasks compared to non-reasoning models of comparable scale. For instance, the reasoning variant of o3 attained an accuracy of 0.7254, exceeding the scores of similarly sized models without such enhancements (Figure 2). This underscores the significant benefit derived from incorporating reasoning-focused mechanisms into LLMs for specialized domains. More details in §4.2.4.
4.2.2 Challenging Nature of ESGenius
Our comprehensive evaluation, detailed in Table 2, reveals that the ESGenius-QA dataset is highly challenging for a broad range of contemporary LLMs. This includes proprietary APIs such as GPT-4o (zero-shot: 0.6364), DeepSeek-R1 (zero-shot: 0.6629), DeepSeek-V3 (zero-shot: 0.6532), and Qwen2.5-Max (zero-shot: 0.6444), as well as leading open-source families like Meta Llama (e.g., Meta-Llama-3-8B zero-shot: 0.6461), Google Gemma (e.g., Gemma-3-12B-Instruct zero-shot:0.6514, Qwen2.5-14B zero-shot: 0.6734). The best zero-shot performance in the table is 0.7254 from the proprietary o3 model, while the highest RAG-enhanced score is 0.8336 with the open-source Gemma-3-27B-Instruct model. This gap underscores both the discriminative nature of ESGenius and the benchmark’s heightened ESG reasoning demands.
4.2.3 RAG Performance
Integrating ESG-specific knowledge through long-context RAG frequently yields substantial performance gains over zero-shot capabilities. For instance, the DeepSeek-R1-Distill-Qwen series demonstrates significant improvements across model sizes: the 1.5B model improves by 37.36% (0.3134 to 0.4305), the 7B model by 29.63% (0.5018 to 0.6505), and the 14B model by 26.07% (0.6382 to 0.8046). Other notable examples include Gemma-3-27B, which achieves a remarkable 141.57% improvement (0.2165 to 0.5229), and QwQ-32B with a substantial 95.23% increase (0.3900 to 0.7614). Similarly, Qwen3-4B improves by 27.75% (0.6188 to 0.7905), and the Qwen2.5-1.5B-Instruct model increases by 27.13% (0.5484 to 0.6972). However, RAG does not universally enhance performance. Some models actually perform worse under the RAG configuration in this study—notably, Qwen2.5-0.5B shows a slight decline of 1.14% (0.5458 to 0.5396), while its instruction-tuned counterpart experiences a more significant 7.48% decrease (0.5775 to 0.5343). These results suggest that for some smaller architectures, additional context may function as noise or exceed their ability to leverage long sequences effectively. Overall, the effectiveness of RAG appears strongly dependent on model architecture and intrinsic capacity, yet it often allows many smaller models to outperform the zero-shot performance of much larger ones.
4.2.4 Reasoning Models Analysis
Table 2 flags “Reasoning Focus” (Rea: Yes) models. Those explicitly oriented or fine-tuned for reasoning, such as DeepSeek-R1-Distill-Qwen and DeepSeek-R1-Distill-Llama (Open Source) and proprietary offerings like o3 (300B, Rea: Yes, zero-shot: 0.7254) or o4-mini (3B, Rea: Yes, zero-shot: 0.6945), frequently excel in zero-shot mode. For example, the reasoning-oriented DeepSeek-R1-Distill-Qwen series shows strong RAG performance and gains (14B model: 0.8046, 1.5B model improves by 37.36%), while QwQ-32B (Rea: Yes) also achieves a 95.23% improvement under RAG. Comparisons with non-reasoning peers reinforce this pattern. For instance, o4-mini (3B, Rea: Yes) reaches 0.6945 zero-shot, well above non-reasoning 3B models such as Llama-3.2-3B (zero-shot: 0.6074) or Qwen2.5-3B (zero-shot: 0.5889). These observations suggest that training or architectural choices emphasizing reasoning bolster ESG question-answering, especially for multi-step logical tasks.
4.2.5 Instruct-Tuned Models Analysis
Our analysis of instruction-tuned models (marked “I-T: Yes” in Table 2) reveals varied performance in zero-shot settings. Several models demonstrate clear benefits from instruction tuning: Google’s Gemma-3-1B improves substantially from 0.2421 to 0.5000, while Gemma-3-27B shows an even more dramatic increase from 0.2165 to 0.6356. Similar positive trends appear in Alibaba’s Qwen2.5-0.5B (0.5458 to 0.5775) and Meta’s Llama-3.2-1B (0.3609 to 0.5986). However, instruction tuning can also lead to performance degradation in some cases: Meta-Llama-3.1-8B declines from 0.6699 to 0.6382, Qwen2.5-1.5B drops from 0.6391 to 0.5484, and Qwen2.5-14B decreases from 0.6734 to 0.6197. The impact of instruction tuning becomes particularly interesting when combined with RAG. In this context, instruction-tuned models frequently demonstrate superior improvement margins compared to their standard counterparts. For instance, the instruction-tuned version of Gemma-3-12B achieves 0.8380 (28.64% improvement) while its non-instruction-tuned variant reaches only 0.6857 (5.26% improvement). Similarly, the instruction-tuned Llama-3.1-8B attains 0.7993 (25.24% improvement) compared to the standard model’s 0.7650 (14.20% improvement).
4.2.6 Error Analysis
Sorting the heatmap (§4.2.8) columns by failure rate exposes a long tail of adversarial-grade questions that no model answered correctly. Appendix C.1 reproduces three such zero-accuracy cases: Question 432 from the SDG 2024 Report probes subtle regional disparities in per-capita renewable electricity capacity; Question 635, drawn from IPCC AR6 WG III, asks for the most direct barrier to ESCO adoption, where models consistently misattributed the bottleneck to regulation or awareness rather than the report’s emphasis on asymmetric information and split incentives; and Question 1006, based on the SASB Chemicals standard, tests the definition of a recordable workplace incident. Across these and similar examples, several error patterns merit closer investigation: a reliance on surface keyword overlap between prompt and distractor, and a tendency to over-generalize to well-known policy frameworks when questions hinge on narrower ESG standards, suggesting that more fine-grained analyses are needed. Future work could therefore combine quantitative error clustering with qualitative case studies, cross-model comparisons, and expert-in-the-loop probing to develop a more comprehensive understanding of the root causes of such failures.
4.2.7 Sub-topic Analysis
A finer-grained sub-topic analysis provides additional insight into model strengths and weaknesses across ESG dimensions. While the current benchmark draws from seven ESG source collections to ensure broad coverage, disaggregating results by ESG pillar (Environmental, Social, Governance) and their sub-themes (§3.1) offers more actionable perspectives. In practice, Environmental spans climate change, biodiversity, pollution control, circular economy, energy transition, etc; Social covers labor rights, diversity and inclusion, community impacts, human health and safety, equity in supply chains, etc; Governance addresses board accountability, anti-corruption, transparency, reporting compliance, risk management, etc; and Cross-Cutting themes such as just transition, Indigenous knowledge, digital governance further blurring boundaries, etc. This infrastructure lays the foundation for mapping each question to its corresponding ESG pillar or sub-theme. In future versions, we will expand this perspective by (i) reporting per-theme performance scores and (ii) discussing how this lens can guide targeted dataset expansion. As part of this process, we have begun annotating each question with an E, S, G, or Cross-subtopic label. For example, Question ID 1—“According to the IPCC AR6 Synthesis Report, which statement would most likely lead to increased vulnerability of biodiversity and ecosystem services while simultaneously reducing carbon uptake, based on the relationships described?”—is labeled E (Environmental, Biodiversity). Question ID 578—“Which initiative best promotes workplace fairness and equal opportunities across different demographic groups within an organization?”—is labeled S (Social, Diversity and Inclusion). In contrast, Question ID 1127—“When an organization determines that indirect economic impacts are a material topic, which of the following best describes the conditions under which it can omit specific disclosures without violating reporting requirements, according to GRI 203: Indirect Economic Impacts 2016?”—is labeled G (Governance, Reporting Compliance). This annotation process is ongoing and will inform future analytical releases, enabling deeper performance breakdowns across ESG pillars and sub-themes.
4.2.8 Interactive Visualization
Figure 9 presents an interactive heatmap visualization of model performance across 50 LLMs (ranging from 0.5 to 671 billion parameters) and all 1,136 MCQs in ESGenius. The visualization arranges models (rows) by their overall zero-shot accuracy and questions (columns) by difficulty, creating a clear left-to-right progression from predominantly red cells (incorrect/unanswered questions) to green cells (correct answers). Interactive features enhance analysis - hovering over any cell reveals detailed information including the model name and rank, question ID, difficulty score, complete prompt, answer options, ground truth, and the model’s prediction. An example of this interactive tooltip is shown in Figure 10. This dynamic visualization transforms raw accuracy data into an intuitive diagnostic tool for analyzing model performance patterns and systematic weaknesses. The interactive visualization is publicly accessible at https://angel-ntu.github.io/ESGenius.
5 Conclusion
We presented ESGenius, the first comprehensive QA benchmark dedicated to probing LLMs on the breadth and depth of ESG and sustainability knowledge. It unifies two tightly integrated components: ESGenius-QA, a large-scale, expert-validated MCQs dataset, and the ESGenius-Corpus, a curated and extensible collection of authoritative ESG sources that anchors every question with verifiable evidence. Our two-stage evaluation protocol—Zero-Shot followed by RAG—shows that grounding responses in curated evidence boosts accuracy by 15–30 percentage points, with smaller models enhanced by RAG often surpassing much larger zero-shot models. This demonstrates that domain integration and transparent grounding are more decisive than sheer scale alone. By open-sourcing data, code, and a live and interactive leaderboard, ESGenius establishes a reproducible and continually updated yardstick for sustainability-aware AI. Our subsequent work will expand this mission to multi-modal ESG tasks Zhang et al. (2025), alongside expert-in-the-loop auditing, paving the way toward trustworthy ESG decision support.
6 Limitations
While ESGenius aspires to provide a holistic evaluation of ESG understanding, several caveats must be acknowledged.
Coverage trade-offs.
The benchmark currently includes only seven sources, which, while representing key frameworks and standards, cannot comprehensively encompass all ESG topics, industry-specific guidelines, regional regulations, or emerging sustainability frameworks.
Expert dependency and scalability.
The creation of high-fidelity questions and the validation of answers are heavily dependent on scarce domain experts, making the process capacity-limited and prone to disciplinary bias. In future iterations, we plan to explore scalable strategies such as expert-in-the-loop active sampling or agreement-based adjudication to sustain long-term benchmark growth.
Format constraints.
The reliance on standardized MCQs, although enabling large-scale scoring, compresses the nuanced reasoning and synthesis required in real-world ESG analysis and may therefore fall short of fully reflecting the complexity of sustainability decision-making.
Standard volatility.
Frameworks such as ISSB guidance and IPCC assessments evolve rapidly, necessitating continual updates, corpus refreshes, and careful version control to maintain relevance.
Language limitations.
The current focus on English-only corpora risks under-representing non-English ESG documents, local regulations, and region-specific sustainability practices. Expanding coverage to multilingual corpora is a key priority for improving global relevance.
Exclusion of visual elements.
ESG reports often include key visual disclosures—charts, graphs, maps, and time-series plots. Our text-only design excludes multimodal reasoning. While text remains central, future versions will integrate visuals via OCR pipelines and vision–language models to better capture ESG reporting.
Metrics and evaluation scope.
Reliance on binary accuracy misses partial correctness and deeper ESG reasoning. We plan to add metrics such as explanation consistency, factual groundedness, and chain-of-thought plausibility. The MCQ format, while limited, offers scalability and comparability, as seen in MMLU and MMMU. We are exploring richer formats—open-ended answers, reasoning traces, and retrieval tasks—for better alignment with real-world ESG analysis.
Bias in LLM-generated questions.
LLMs are widely used in benchmarks for question generation and evaluation. Our pipeline adds expert verification to ensure factuality and relevance, but both models and humans can introduce bias. While we emphasize consistency and oversight, residual biases remain. We welcome validated alternatives to further reduce bias in future versions.
Limitations of the RAG implementation.
We use a simple RAG setup for reproducibility, deployment with smaller LLMs, and clear attribution of performance gains to the dataset and task design. Despite its simplicity, it already yields strong improvements. Future versions will integrate more advanced retrieval to enhance robustness and real-world alignment.
Human baseline performance.
Without human benchmarks, model scores lack context. We are developing a crowdsourcing interface for experts to provide evaluations, establishing baselines that will ground interpretation of results and guide future benchmarking.
Copyright limits.
Licensing restrictions constrain the inclusion of certain proprietary ESG standards, limiting exhaustive coverage and potentially omitting significant industry-specific frameworks.
7 Ethical Considerations
Purpose and scope.
ESGenius is released to advance trustworthy, evidence–grounded research on LLMs Environmental, Social, and Governance (ESG) question answering.
Sources, copyright, and licensing.
The ESGenius-Corpus aggregates authoritative but heterogeneous materials (e.g., GRI, SASB/ISSB, IPCC, TCFD, CDP, SDGs in §3.2). Many of these works are copyrighted and/or distributed under specific use terms. We therefore: (i) include only publicly available documents or clearly reference official distribution pages; (ii) link every question to minimal excerpts sufficient for verification, avoiding wholesale redistribution; and (iii) ship documentation that records provenance and access routes for each source. Users of ESGenius must review and comply with the licenses and terms of the original sources; the benchmark is provided strictly for academic research and reproducibility.
LLM assistance and data leakage.
LLMs were used only to propose candidate MCQs; six domain experts double–validated all retained questions and three senior practitioners consulted the process (§3.1). To reduce circularity, we (i) enforced strict evidence requirements per question, (ii) rejected or revised questions that were not unambiguously supported, and (iii) evaluated models—including the assisting model—via API in an isolated session, with no access to construction traces. Each question is mapped to source passages (§3.1), and a “Not sure (Z)” option was added to discourage hallucination and reward calibrated abstention (§3.1). Beyond content generation, LLMs were occasionally employed as supportive tools for surface-level polishing, such as refining phrasing in writing or improving code readability and consistency. These uses were strictly non-substantive and did not influence the design, validation, or evaluation of ESGenius itself.
Annotation practice and labor.
All validators were experienced in sustainability or NLP and gave informed consent to participate. The workflow emphasized factual verification, clarity, and distractor plausibility, with systematic rejection of ambiguous questions (§3.1). We credit contributors as co–authors, and we encourage future adopters to follow fair labor practices when extending the dataset.
Fairness, representation, and epistemic balance.
ESGenius currently emphasizes English–language, investor–oriented frameworks and global standards (e.g., IFRS S1/S2, SASB) over national, local, Indigenous, or multilingual sources (§6). This may reproduce an institutional or Global–North lens and under–represent region–specific norms or community priorities. We flag this as an ethical limitation and design the corpus to be modular and extensible3.2). Users should avoid over–generalizing conclusions beyond the covered sources and geographies.
Dual–use risks and misuse.
Benchmarking can be gamed or misapplied. Examples include (i) training to the test (overfitting to known questions) and (ii) using ESG QA capability claims to enable persuasive but ungrounded communication or greenwashing. To mitigate these risks we: (a) link every question to authoritative evidence and encourage evaluation settings that require citations (RAG, §4.1.2); (b) provide versioning and clear changelogs; (c) retain the “Z” option to penalize confident guessing; and (d) discourage using ESGenius scores as sole evidence of real–world compliance or as a checklist for circumventing disclosure obligations. We further recommend reporting retrieval quality and evidence coverage alongside accuracy.
Leaderboard and claims.
Model rankings can invite outsized or decontextualized claims. We caution that (i) ESGenius is one lens among many (§6); (ii) zero–shot accuracy reflects parametric knowledge, whereas RAG performance reflects retrieval and grounding design (§4.2); and (iii) differences of a few percentage points may not be practically meaningful without error and subtopic analyses (§4.2.7). Public results should disclose prompts, decoding settings, retrieval parameters, and version identifiers to support comparability and reproducibility.
Safety of model outputs.
ESG advice or interpretations can carry legal, financial, or societal consequences. ESGenius is not a substitute for professional guidance. We encourage deployers to (i) present grounded citations by default; (ii) surface uncertainty via calibrated abstention; and (iii) keep a expert-in-the-loop for consequential uses (e.g., risk management, regulatory filings). Our findings (e.g., stronger gains with RAG) underscore that evidence-linked responses are ethically preferable to unsupported generations.
Environmental impact.
Benchmarking LLMs consumes energy and incurs carbon costs. Our experiments used a single DGX A100 node (4 $\times$ 80GB) to limit overhead. We encourage (i) reporting hardware and token budgets, (ii) batching and caching retrieval, and (iii) preferring smaller models with strong RAG performance when feasible (§4.2.1).
Responsible release and governance.
We release ESGenius with (i) documented data lineage and versioning; (ii) explicit usage guidelines discouraging compliance automation or deceptive communication; and (iii) maintenance plans for updates as standards evolve (§3.2, §5). Community feedback on coverage gaps, bias, and error reports is welcome, and validated contributions will be incorporated through transparent review.
Acknowledgments
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).
Appendix
Appendix A ESGenius-QA
A.1 Automated Preparation of Candidate MCQs
This appendix details the fully-automated preparation stage that converts raw knowledge sources into a pool of candidate MCQs. (The complete algorithm is detailed in Algorithm 1.) These candidate questions subsequently undergo expert review, editing, and validation before potential inclusion in ESGenius-QA. The goals of the preparation stage are twofold: (i) breadth—to cover as many distinct passages as possible from the source corpus, and (ii) difficulty—to generate questions that force LLMs to reason beyond surface-level facts.
Input corpus.
Documents are first collected from authoritative sources in PDF format. Each PDF is then processed independently via the pipeline to enable precise tracking of corpus coverage. For encrypted PDFs, the pipeline attempts in-memory decryption where possible; if decryption fails, the file is skipped with a warning. Throughout processing, the system maintains detailed logs of extracted page counts and flags any instances of unsupported encryption formats.
Text extraction and chunking process.
Using PyPDF2 Fenniak et al. (2022), every page’s text is cleaned (whitespace collapse and control-character stripping) and stored in a page map. A chunk selector then samples passages that satisfy length constraints
$150\leq|{\rm chunk}|\leq 4500\text{ characters},$
optionally concatenating the next page when the combined length remains under the upper bound. Each chunk is tagged with its page span (e.g., “12–13”) so that provenance is preserved.
Prompted multiple-choice question generation.
For every selected chunk the pipeline sends a carefully engineered prompt (§A) to the Qwen Max’s DashScope endpoint111qwen-max-2025-01-25, temperature 0, deterministic for reproducibility. that requests exactly one “extremely challenging” MCQ. Mandatory instructions enforce (i) deep reasoning across multiple sentences, (ii) near-miss distractors, (iii) answerability from the excerpt alone, and (iv) strict JSON output. A fixed seed (42) controls chunk sampling so that runs are reproducible.
Structural validation and metadata augmentation.
The returned JSON is sanitised, parsed, and validated. Missing or malformed keys trigger rejection. The correct answer text is cross-checked with the declared key, and a fallback correction is applied if necessary. A universal Z option (“Not sure”) is appended to every question to enable abstention analysis. Each validated MCQ is wrapped in a metadata envelope containing {id, source_pdf, page_range, difficulty, generation_timestamp}.
Iterative difficulty refinement process.
Because the pipeline aims for a target zero-shot accuracy of $\leq 50\%$ on an external evaluator (we use Qwen Max as the evaluator), it performs up to 15 iterative refinement rounds to achieve this threshold:
  1. 1.
    Test the current pool; label each MCQ with the model’s chosen answer and correctness.
  2. 2.
    Remove questions that the evaluator answers correctly.
  3. 3.
    Replace the removed questions by sampling new chunks and repeating the generation step, maintaining the original pool size.
  4. 4.
    Abort early if (a) accuracy falls below the target, (b) authentication errors reach the threshold (default 5), or (c) no new valid chunks remain.
Consistency checking and question deduplication.
Beginning with round 2 the pipeline checks that an MCQ’s evaluator result is stable across rounds; inconsistent questions are discarded. After refinement, a pairwise LLM-based similarity check removes near-duplicate questions that probe the same underlying concept. Only the first occurrence is retained.
Output artifacts generation and storage.
Two JSON files are written atomically after every processed PDF: all_candidates_untested.json stores every syntactically valid question, whereas final_round_tested.json retains only the post-refinement pool together with evaluator metadata. Incremental saves ensure that partial progress survives crashes, and IDs are globally unique across sessions.
Transition to expert review process.
The resulting candidate pool is then passed to the expert review process detailed in §3.1. This marks the transition from automated preparation to rigorous human validation and refinement of each question.
Exact Prompt Template
Context: Based *only* on the following text excerpt from page(s) {page_range} of the document ’{doc_name}’:
{text_excerpt}
Task: Generate ONE **extremely challenging** multiple-choice question (MCQ)
with a target difficulty level of ’{difficulty}’ specifically designed to be difficult for advanced LLMs like Qwen-Max.
The question must rigorously test deep comprehension, critical reasoning, inference, and synthesis *strictly* based on
the provided text excerpt. Avoid simple factual recall.
**Mandatory Instructions for Difficulty:**
1.  **Question Focus:** The question must target subtle details, implicit relationships, logical consequences,
nuanced comparisons, or the underlying purpose/function of specific information within the text. It *must* require
synthesizing information from *multiple distinct sentences or points* within the excerpt.
2.  **Distractor Design (CRITICAL):** Create highly plausible but definitively incorrect distractors. These distractors
should:
    * Be semantically very close to the correct answer or other concepts in the text.
    * Represent common misconceptions or logical fallacies related to the topic, *but be directly contradicted by
    the provided text*.
    * Exploit potential ambiguities or require careful reading to identify subtle inaccuracies.
    * Avoid being trivially wrong or easily eliminated. All options (A, B, C, D) should seem plausible on a
    superficial reading.
3.  **Avoid External Knowledge:** The question MUST be unambiguously answerable *solely* from the information
contained within the provided text excerpt. The difficulty must come from the complexity of reasoning *within* the text,
not from needing outside information.
4.  **Direct Questioning:** The text of the generated ‘question‘ itself must **not** contain phrases like
"Based on the text excerpt", "According to the document...", "In the provided context...", etc. Frame the question
directly about the content.
5.  **Advanced Strategies:** Employ techniques like identifying cause-effect relationships only implied,
analyzing the necessity or sufficiency of conditions mentioned, or evaluating the scope/limitations
of claims made in the text. Carefully consider using negation/exception questions ("Which of the following is NOT...")
only if they significantly increase cognitive load beyond simple recall.
**Self-Correction Check (Internal Thought Process before outputting):** Does this question *genuinely* require
multi-step reasoning or deep inference based *only* on the text? Are the distractors *truly* deceptive and require
careful analysis of the excerpt to disprove? Is it significantly harder than a simple keyword search
or basic comprehension question?
Output Format: Respond ONLY with a single, valid JSON object matching this exact structure:
  "question": "String: The extremely challenging question text (must not refer to the prompt or text source).",
  "options": { "A": "String: Highly plausible but incorrect
  Option A.", "B": "String: Highly plausible but incorrect Option B.", "C": "String: The single, unambiguously correct
  answer based *only* on the text.", "D": "String: Highly plausible but incorrect Option D." },
  "correct_answer_text": "String: The exact text of the correct answer option (must match one of the options exactly).",
  "correct_option_key": "String: The key (’A’, ’B’, or ’C’, or ’D’) corresponding to the correct answer."
**Important:** Ensure the ‘correct_option_key‘ points to the option that IS the correct answer. Randomize the position
of the correct answer (A, B, C, or D). Do not include any introductory text, explanations, or markdown formatting like
‘‘‘json ... ‘‘‘ before or after the JSON object. Output *only* the JSON.
1
2Hyper-parameters:
3 Chunk length $L\in[150,4500]$ chars;  LLM = qwen-max-2025-01-25, T=0T{=}0;  seed =42=42;  max rounds $R_{\max}=15$;  target acc. $\leq 50\%$;  auth-error cap $E_{\max}=5$.
4
Input : PDF folder knowledge_source
Output : all_candidates_untested.json, final_round_tested.json
5
6foreach PDF $d$ in folder do // independent handling
7   decryptIfPossible(d)(d)
8 if unsupported encryption then continue
9
  pages $\leftarrow$ ExtractPages(d)
 // cleaned text
10   chunks $\leftarrow$ SelectChunks(pages,$L$)
11
12 foreach chunk $c$ in chunks do
13    $\mathit{mcq}\leftarrow$ PromptLLM(c)(c)
[…]
15         append option $\mathbf{Z}$ (“Not sure”)
16         add metadata $\{\texttt{id},d,\texttt{page\_range},\texttt{difficulty},\texttt{timestamp}\}$
17         save to all_candidates_untested.json
18
19      end if
20
21   end foreach
22
23  round $\leftarrow 1$;
24   acc $\leftarrow 1$;
25   errs $\leftarrow 0$
26 while $round\leq R_{\max}$ and $acc>0.5$ do
27      acc $\leftarrow$ EvaluatePool()
28      remove correctly answered MCQs
29    if $errs\geq E_{\max}$ or no new chunks then break
[…]
31    $round\leftarrow round+1$
32
33   end while
34
35  drop unstable MCQs across rounds
36   deduplicate with LLM similarity check (keep first)
37   atomically write surviving pool $\rightarrow$ final_round_tested.json
38
39 end foreach
40
41Expert review: domain specialists verify facts, polish wording, rebalance difficulty, and accept/reject questions for release.
42
Algorithm 1 Automated pipeline for generating candidate MCQs
Metric Value
Questions
Entries/Tokens/Vocab 1 136 / 40 983 / 2 896
Mean (Median) 36.08 (35)
Q1 – Q3 30 – 40
Range 18 – 94
Options
Entries/Tokens/Vocab 5 680 / 88 304 / 6 810
Mean (Median) 15.55 (17)
Q1 – Q3 12 – 21
Range 1 – 54
Source Text
Entries/Tokens/Vocab 1 136 / 550 200 / 18 826
Mean (Median) 484.33 (467)
Q1 – Q3 390 – 586
Range 33 – 984
Table 1: Token–level profile of the ESGenius-QA. Source texts provide dense evidence (median 467 tokens) supporting higher-order reasoning.
A.2 LLM’s Raw Response Validation
Given an LLM’s raw response, the routine first guards against None or empty inputs, then accepts “Z” directly as the special unsure option. For other strings it proceeds in two stages:
[…]
Robust cleanup: Otherwise the string is sanitized: (i) replace all non-alphabetic symbols by spaces, (ii) drop any word containing lowercase letters, (iii) retain only words composed of A--D,Z. The result is deduplicated and sorted as above. If the cleaned string is empty, contains a lone “Z”, or mixes “Z” with other letters, the function outputs INVALID_ANSWER_MARKER.
This design accepts the minimal valid alphabet while aggressively filtering free-form text, punctuation, and lower-case distractors that often appear in LLM generations, to guarantee that the response is a valid multiple-choice answer.
A.3 ESGenius-QA Example Question Structures
A.3.1 Zero-Shot Evaluation Example
Question ID: 1
Question. According to the IPCC AR6 Synthesis Report, which statement would most likely increase the vulnerability of biodiversity and ecosystem services while simultaneously reducing carbon uptake, given the relationships described?
A. Implementing targeted management strategies for ocean ecosystems without addressing global warming.
B. Failing to rebuild overexploited fisheries while achieving significant reductions in fossil-fuel dependency.
C. Limiting global warming but neglecting land-restoration efforts and cooperative decision-making with Indigenous Peoples.
D. Prioritizing disaster-risk management and early-warning systems without integrating ecosystem considerations into climate policies.
Z. Not sure
Correct Answer: A
A.3.2 RAG Evaluation Example
Question ID: 580
Reference. Page 213, SUS Report.pdf
Question. Which of the following conclusions about Fiji’s progress toward sustainable development can be most reasonably inferred from the data trends and gaps presented in the Sustainable Development Report 2024?
A. Fiji has achieved near-universal access to clean water but faces significant challenges in reducing urban slum populations.
B. Fiji’s environmental sustainability efforts are hindered primarily by high deforestation rates and low biodiversity protection.
C. Fiji excels in reducing income inequality, as evidenced by a low Gini coefficient and minimal reliance on imports for nitrogen emissions.
D. Fiji demonstrates strong performance in internet accessibility but shows untracked research and development expenditure.
Z. Not sure
Correct Answer: D
Retrieved Source Text. Performance by Indicator5. Country Profiles Sustainable Development Report 2024 – The SDGs and the UN Summit of the Future 201
* Imputed data point, ** Not applicable NA = Data not available
FIJI
SDG 9 – Industry, Innovation and Infrastructure
Rural population with access to all-season roads (%) 96.7 (2024) ••
Population using the internet (%) 85.2 (2022) •A
Mobile broadband subscriptions (per 100 population) 76.1 (2021) •A
Logistics Performance Index: Infrastructure score (worst 1–5 best) 2.2 (2023) •G
The Times Higher Education Universities Ranking: Average score of top 3 universities (worst 0–100 best) 30.5 (2024) ••
Articles published in academic journals (per 1 000 population) 0.6 (2022) •A
Expenditure on research and development (% of GDP) NA NA ••
SDG 10 – Reduced Inequalities
Gini coefficient 30.7 (2019) ••
Palma ratio 1.1 (2019) ••
SDG 11 – Sustainable Cities and Communities
Proportion of urban population living in slums (%) 9.4 (2020) •D
Annual mean concentration of $PM_{2.5}$ (µg/m³) 7.4 (2022) •D
Access to improved water source, piped (% of urban population) 98.4 (2022) •A
Population with convenient access to public transport in cities (%) 19.2 (2020) ••
SDG 12 – Responsible Consumption and Production
Municipal solid waste (kg/capita/day) 0.6 (2011) ••
Electronic waste (kg/capita) 6.1 (2019) ••
Production-based air pollution (DALYs per 1 000 population) NA NA ••
Air pollution associated with imports (DALYs per 1 000 population) NA NA ••
Production-based nitrogen emissions (kg/capita) NA NA ••
Nitrogen emissions associated with imports (kg/capita) NA NA ••
Exports of plastic waste (kg/capita) 0.6 (2022) •A
SDG 13 – Climate Action
$CO_{2}$
emissions from fossil-fuel combustion and cement production (t $CO_{2}$/capita) 1.2 (2022) •A
GHG emissions embodied in imports (t $CO_{2}$/capita) NA NA ••
$CO_{2}$ emissions embodied in fossil-fuel exports (kg/capita) 0.0 (2022) ••
SDG 14 – Life Below Water
Mean area that is protected in marine sites important to biodiversity (%) 16.5 (2023) •D
Ocean Health Index: Clean Waters score (worst 0–100 best) 74.1 (2023) •D
Fish caught from overexploited or collapsed stocks (% of total catch) 9.0 (2018) •A
Fish caught by trawling or dredging (%) 0.0 (2019) ••
Fish caught that are then discarded (%) 7.3 (2019) •D
Marine biodiversity threats embodied in imports (per million population) 0.3 (2018) ••
SDG 15 – Life on Land
Mean area that is protected in terrestrial sites important to biodiversity (%) 11.2 (2023) •D
Mean area that is protected in freshwater sites important to biodiversity (%) 0.1 (2023) •D
Red List Index of species survival (worst 0–1 best) 0.69 (2024) •G
Permanent deforestation (% of forest area, 3-year average) 0.1 (2022) •A
Imported deforestation (m²/capita) NA NA ••
SDG 16 – Peace, Justice and Strong Institutions
Homicides (per 100 000 population) 2.2 (2020) •D
Crime is effectively controlled (worst 0–1 best) NA NA ••
Unsentenced detainees (% of prison population) 19.9 (2021) •A
Birth registrations with civil authority (% of children under 5) 86.6 (2021) ••
Corruption Perceptions Index (worst 0–1 best) 52.0 (2023) ••
Children involved in child labor (%) 16.7 (2021) ••
Exports of major conventional weapons (TIV constant million USD per 100 000 population) * 0.0 (2023) ••
Press Freedom Index (worst 0–1 best) 71.2 (2024) •A
Access to and affordability of justice (worst 0–1 best) NA NA ••
Timeliness of administrative proceedings (worst 0–1 best) NA NA ••
Expropriations are lawful and adequately compensated (worst 0–1 best) NA NA ••
SDG 17 – Partnerships for the Goals
Government spending on health and education (% of GDP) 9.3 (2021) •A
Government revenue excluding grants (% of GDP) 19.0 (2021) •G
Corporate Tax Haven score (best 0–100 worst) * 0 (2021) ••
Statistical Performance Index (worst 0–100 best) 63.2 (2022) •S
Index of countries’ support to UN-based multilateralism (worst 0–100 best) 88.3 (2023) ••
SDG 1 – No Poverty
Poverty headcount ratio at \$2.15/day (2017 PPP, %) 1.6 (2024) •D
Poverty headcount ratio at \$3.65/day (2017 PPP, %) 7.3 (2024) •D
SDG 2 – Zero Hunger
Prevalence of undernourishment (%) 6.6 (2021) •A
Prevalence of stunting in children under 5 years of age (%) 7.2 (2021) ••
Prevalence of wasting in children under 5 years of age (%) 4.6 (2021) ••
Prevalence of obesity, BMI $\geq$ 30 (% of adult population) 33.8 (2022) •G
Human Trophic Level (best 2–3 worst) 2.2 (2021) •D
Cereal yield (tonnes per hectare of harvested land) 4.1 (202…)
Appendix B Prompt Template
For MCQs (4 options + 1 Not sure with 1 answer), we use the following prompt templates:
B.1 Zero-Shot Prompt
You are an expert in ESG (Environmental, Social, Governance) and Sustainability related topics. Answer the question with a single letter based on authoritative knowledge. Each option content is case-sensitive.
Question: [Question text]
Options:
A: [Option A text]
B: [Option B text]
C: [Option C text]
D: [Option D text]
Z: Not sure
Answer: <Model’s response goes here>
B.2 RAG Prompt
Context: [source text]
You are an expert in ESG (Environmental, Social, Governance) and Sustainability related topics. Answer the question with a single letter based on authoritative knowledge. Each option content is case-sensitive.
Question: [Question text]
Options:
A: [Option A text]
B: [Option B text]
C: [Option C text]
D: [Option D text]
Z: Not sure
Answer: <Model’s response goes here>
Appendix C Main Experimental Results Table
Type Family Model Size S.G. I-T Rea Zero-Shot RAG Improvement
Open Source DeepSeek DeepSeek-R1-Distill-Qwen 1.5B M No Yes 0.3134 0.4305 37.36%
DeepSeek-R1-Distill-Qwen 7B L No Yes 0.5018 0.6505 29.63%
DeepSeek-R1-Distill-Qwen 14B L No Yes 0.6382 0.8046 26.07%
DeepSeek-R1-Distill-Qwen 32B XL No Yes 0.6312 0.8143 29.01%
DeepSeek-R1-Distill-Llama 8B L No Yes 0.5502 0.6250 13.60%
DeepSeek-R1-Distill-Llama 70B XL No Yes 0.6505 0.8170 25.60%
Google Gemma Gemma-3 Team et al. (2025) 1B M No No 0.2421 0.2526 4.33%
Gemma-3 1B M Yes No 0.5000 0.5977 19.54%
Gemma-3 4B M No No 0.6083 0.6860 12.77%
Gemma-3 4B M Yes No 0.6144 0.7518 22.36%
Gemma-3 12B L No No 0.6514 0.6857 5.26%
Gemma-3 12B L Yes No 0.6514 0.8380 28.64%
Gemma-3 27B L No No 0.2165 0.5229 141.57%
Gemma-3 27B L Yes No 0.6356 0.8336 31.15%
Meta Llama Meta-Llama-3 Grattafiori et al. (2024) 8B L No No 0.6461 0.7324 13.36%
Llama-3.1 8B L No No 0.6699 0.7650 14.20%
Llama-3.1 8B L Yes No 0.6382 0.7993 25.24%
Llama-3.2 1B M No No 0.3609 0.3680 2.00%
Llama-3.2 1B M Yes No 0.5986 0.6452 7.79%
Llama-3.2 3B M No No 0.6074 0.6831 12.48%
Llama-3.2 3B M Yes No 0.5968 0.7218 20.95%
Llama-3.3 70B XL Yes No 0.6576 0.7887 20.00%
Alibaba Qwen Qwen2.5 Yang et al. (2024) 0.5B S No No 0.5458 0.5396 -1.14%
Qwen2.5 0.5B S Yes No 0.5775 0.5343 -7.48%
Qwen2.5 1.5B M No No 0.6391 0.6928 8.40%
Qwen2.5 1.5B M Yes No 0.5484 0.6972 27.13%
Qwen2.5 3B M No No 0.5889 0.7632 29.60%
Qwen2.5 3B M Yes No 0.5871 0.5211 -11.24%
Qwen2.5 7B L No No 0.6496 0.8055 23.99%
Qwen2.5 7B L Yes No 0.6276 0.7967 27.27%
Qwen2.5 14B L No No 0.6734 0.8231 22.22%
Qwen2.5 14B L Yes No 0.6197 0.8231 32.83%
Qwen2.5 32B XL No No 0.6593 0.8081 22.55%
Qwen2.5 32B XL Yes No 0.6039 0.8247 36.57%
Qwen2.5 72B XL No No 0.6188 0.7201 16.39%
Qwen2.5 72B XL Yes No 0.6347 0.8257 29.78%
Qwen2.5-1M Yang et al. (2025eb) 7B L Yes No 0.6206 0.8063 29.92%
Qwen2.5-1M 14B L Yes No 0.6268 0.8222 28.01%
QwQ Team (2025) 32B XL No Yes 0.3900 0.7614 95.23%
Qwen3 Yang et al. (2025) 0.6B S No No 0.2896 0.0942 -67.47%
Qwen3 1.7B M No No 0.5836 0.6937 18.87%
Qwen3 4B M No No 0.6188 0.7905 27.75%
Qwen3 8B L No No 0.6021 0.6708 11.41%
Proprietary API DeepSeek DeepSeek-R1 Guo et al. (2025) 671B XXL No Yes 0.6629 - -
DeepSeek-V3 Liu et al. (2024) 671B XXL No No 0.6532 - -
Alibaba Qwen Qwen2.5-Max (qwen-max-2025-01-25) A22B†(MoE 325B) XL No Yes 0.6444 - -
OpenAI GPT GPT-4o-mini 8B† L No No 0.6268 - -
GPT-4o Hurst et al. (2024) 200B† XXL No No 0.6364 - -
o4-mini 3B† M No Yes 0.6945 - -
o3 300B† M No Yes 0.7254 - -
Table 2: Comprehensive ESGenius results showing LLM performance under Zero-Shot and RAG settings. S.G. denotes Size Group (S: Small ($\leq$1B), M: Medium (1–7B), L: Large (7–30B), XL: Extra Large (30–100B), XXL: Extra Extra Large (>>100B)); I-T: Instruction Tuning; Rea: Reasoning Focus. Improvement (%): $\frac{\text{RAG}-\text{Zero-Shot}}{\text{Zero-Shot}}\times 100$. “–” indicates scores currently unavailable due to technical constraints. † indicates industry size estimates. A ranking bar chart of zero-shot performance is shown in Figure 3.
Figure 3: Zero-Shot model ranking on ESGenius (1136 questions). The chart shows model accuracies across 50 LLMs, with o3 achieving the highest score of 72.54%.
C.1 3 Examples with Very Low Accuracy Across All Models
Question ID: 432
Question. Which statement accurately reflects the relationship between renewable energy adoption and regional disparities as indicated in The Sustainable Development Goals Report 2024?
Options:
A: Developing countries are projected to surpass developed countries in per capita renewable electricity capacity within the next decade based on current growth rates.
B: The disparity in renewable energy capacity between least developed countries (LDCs) and developing countries is expected to close within 15 years if LDCs maintain their current growth trajectory.
C: The installed renewable electricity capacity in least developed countries (LDCs) grew at a faster compound annual growth rate than in developed countries over the past seven years.
D: Landlocked developing States have achieved a higher per capita renewable electricity capacity than small island developing countries but still lag significantly behind developing countries overall.
Z: Not sure
Correct Answer: D
Accuracy: 0%
Question ID: 635
Question. Which factor is most directly responsible for the limited adoption of ESCO business models despite their potential to mitigate financial risks and provide expertise in energy efficiency projects, according to the Climate Change 2022: Mitigation of Climate Change. Working Group III Contribution to the Sixth Assessment Report of the Intergovernmental Panel on Climate Change?
Options:
A: The absence of stringent regulatory frameworks governing energy codes.
B: The insufficient involvement of the public sector in transportation infrastructure projects.
C: The lack of awareness among financial institutions about energy efficiency benefits.
D: The prevalence of asymmetric information and split incentives within firms.
Z: Not sure
Correct Answer: C
Accuracy: 0%
Question ID: 1006
Question. Which of the following best describes a necessary condition for an injury or illness to be classified as a recordable incident under the entity’s disclosure requirements in the Chemicals – Sustainability Accounting Standard?
Options:
A: The injury or illness must result in at least one day away from work or require medical treatment beyond first aid.
B: The injury or illness must be diagnosed by a licensed healthcare professional, regardless of its impact on the employee’s work status.
C: The injury or illness must occur within the establishment but does not need to involve exposure to harmful substances or heavy machinery.
D: The injury or illness must lead to restricted work, job transfer, or loss of consciousness, even if it is not diagnosed by a physician.
Z: Not sure
Correct Answer: B
Accuracy: 0%
Appendix D ESGenius-Corpus
The corpus integrates authoritative frameworks, corporate questionnaires, and scientific assessments that comprehensively cover environmental, social, and governance (ESG) considerations. A detailed quantitative overview of the collection is presented across multiple tables: Table 3 summarizes IPCC climate-science assessments, Table 4 covers GRI Standards, Tables 56 detail SASB’s industry-specific disclosures, Tables 78 outline IFRS/ISSB standards and guidance, Table 9 catalogs TCFD materials, Table 10 documents CDP questionnaires, and Table 11 captures UN Sustainable Development Goals (SDGs) content. The distribution of pages and questions across these sources is visualized in Figures 7 and 8, respectively. From this comprehensive dataset, several key insights emerge:
  1. 1.
    Breadth versus depth. Altogether, these seven sources comprise 231 distinct documents spanning approximately 19,600 pages. While bodies such as SASB and IFRS/ISSB each contribute dozens of relatively concise files, the Intergovernmental Panel on Climate Change (IPCC) anchors the corpus with just seven reports that total over 10,000 pages.
  2. 2.
    Standards-driven core. Internationally recognized ESG standards and disclosure frameworks—GRI, SASB, IFRS/ISSB, TCFD, and CDP—form the backbone of ESGenius. They reflect the most widely adopted practices for sustainability reporting, management, and climate-risk disclosure, offering a robust foundation for both quantitative and qualitative benchmarking.
  3. 3.
    Disclosure and reporting frameworks. Market-facing reporting initiatives (SASB, GRI, TCFD, CDP) collectively emphasize implementation guides and sector- or topic-specific questionnaires, resulting in many individual questions but fewer pages per document. This modular structure facilitates domain-specific data collection and comparability across diverse industries.
  4. 4.
    Science-heavy climate assessments. IPCC assessment reports present the opposite pattern: fewer individual documents but extremely high page counts. This science-heavy text ensures deep coverage of climate-change fundamentals, impacts, and mitigation pathways—an essential knowledge base informing the standards and regulations in the broader ESG ecosystem.
  5. 5.
    Sustainable development anchor. The UN’s SDGs underpin cross-sector and cross-country sustainability objectives. Although comparatively compact in page count, these seminal UN publications situate corporate ESG strategies within the global 2030 Agenda, ensuring broader alignment with international development priorities.
  6. 6.
    Imbalanced density highlights practical challenges. Marked disparities between the distribution of documents and the distribution of pages underscore the varied scope of ESG sources: some (e.g., IPCC) are exhaustive scientific compendiums, while others (e.g., GRI, SASB, IFRS) comprise slimmer but more numerous reference standards. For researchers and practitioners alike, tasks ranging from large-scale summarization to specialized technical queries must navigate this imbalance of question count versus depth.
Taken together, these characteristics demonstrate that ESGenius provides both the breadth (multiple standards, guidance, and scientific anchors) and the depth (tens of thousands of pages) required for evaluating advanced language models on ESG-focused reasoning, retrieval, and generation tasks. Future expansions will incorporate emerging disclosure rules and further national or sectoral guidelines, maintaining the corpus’s comprehensive coverage over time.
Original Document Title Year Size Pages No. Qs
Climate Change 2023 — Synthesis Report IPCC (2023cb) 2023 4.9 MB 186 18
Climate Change 2022: Mitigation of Climate Change. Working Group III Contribution to the Sixth Assessment Report of the Intergovernmental Panel on Climate Change IPCC (2022cb) 2022 74.2 MB 2 042 19
Climate Change 2022: Impacts, Adaptation and Vulnerability. Working Group II Contribution to the IPCC Sixth Assessment Report IPCC (2022) 2022 378 MB 3 675 14
Climate Change 2021: The Physical Science Basis. Working Group I Contribution to the IPCC Sixth Assessment Report IPCC (2021) 2021 275 MB 2 409 14
Climate Change and Land: An IPCC Special Report on climate change, desertification, land degradation, sustainable land management, food security, and greenhouse gas fluxes in terrestrial ecosystems IPCC (2019) 2019 28 MB 874 16
The Ocean and Cryosphere in a Changing Climate: A Special Report of the Intergovernmental Panel on Climate Change IPCC (2019cb) 2019 59.4 MB 765 17
Global Warming of 1.5 °C: An IPCC Special Report on the impacts of global warming of 1.5°C above pre-industrial levels and related global greenhouse gas emission pathways, in the context of strengthening the global response to the threat of climate change, sustainable development, and efforts to eradicate poverty IPCC (2018) 2018 65 MB 631 19
Total - 884.5MB 10 582 117
Table 3: Comprehensive metadata for the seven IPCC reports curated in the ESGenius. This collection represents the complete Sixth Assessment Report (AR6) cycle and key Special Reports from 2018-2023. The AR6 materials include the 2023 Synthesis Report and three Working Group contributions covering physical science (WG1), impacts & adaptation (WG2), and mitigation (WG3). Three thematic Special Reports address land use (SRCCL), oceans & ice (SROCC), and 1.5°C warming pathways (SR15). Totaling 10,582 pages and 117 evaluation questions, these authoritative climate science assessments form a crucial knowledge foundation for ESG analysis. Document sizes range from 4.9MB to 378MB. The collection provides comprehensive coverage of climate science, impacts, and policy responses that inform modern ESG frameworks like TCFD, CSRD, and ISSB standards. All documents are sourced directly from IPCC (https://www.ipcc.ch) and represent peer-reviewed, UN-mandated scientific assessments.
Original Document Title Year Size Pages No. Qs
Consolidated Set of the GRI Standards 2021 gri (2021) 2021 19 MB 677 46
GRI 1: Foundation 2021 gri (2021ab) 2021 1.2 MB 39 6
GRI 2: General Disclosures 2021 gri (2021ad) 2021 1.2 MB 58 11
GRI 3: Material Topics 2021 gri (2021ae) 2021 1.1 MB 30 5
GRI 11: Oil and Gas Sector 2021 gri (2021ac) 2021 2.2 MB 93 14
GRI 12: Coal Sector 2022 gri (2022) 2022 2.1 MB 86 16
GRI 13: Agriculture, Aquaculture and Fishing Sectors 2022 gri (2022ab) 2022 2.5 MB 95 17
GRI 14: Mining Sector 2024 gri (2024ab) 2024 2.6 MB 100 15
GRI 101: Biodiversity 2024 gri (2024) 2024 1.3 MB 50 10
GRI 201: Economic Performance 2016 gri (2016a) 2016 862 KB 16 2
GRI 202: Market Presence 2016 gri (2016b) 2016 834 KB 15 1
GRI 203: Indirect Economic Impacts 2016 gri (2016c) 2016 817 KB 11 3
GRI 204: Procurement Practices 2016 gri (2016d) 2016 821 KB 11 3
GRI 205: Anti-corruption 2016 gri (2016e) 2016 855 KB 16 1
GRI 206: Anti-competitive Behavior 2016 gri (2016f) 2016 829 KB 13 1
GRI 207: Tax 2019 gri (2019) 2019 947 KB 21 3
GRI 301: Materials 2016 gri (2016g) 2016 831 KB 13 1
GRI 302: Energy 2016 gri (2016h) 2016 859 KB 19 1
GRI 303: Water and Effluents 2018 gri (2018) 2018 1 MB 28 4
GRI 304: Biodiversity 2016 gri (2016i) 2016 845 KB 15 1
GRI 305: Emissions 2016 gri (2016j) 2016 936 KB 26 4
GRI 306: Effluents and Waste 2016 gri (2016k) 2016 640 KB 15 1
GRI 306: Waste 2020 gri (2020) 2020 1.7 MB 30 1
GRI 308: Supplier Environmental Assessment 2016 gri (2016l) 2016 851 KB 14 1
GRI 401: Employment 2016 gri (2016m) 2016 860 KB 16 2
GRI 402: Labor/Management Relations 2016 gri (2016n) 2016 855 KB 13 1
GRI 403: Occupational Health and Safety 2018 gri (2018ab) 2018 1.1 MB 35 6
GRI 404: Training and Education 2016 gri (2016o) 2016 837 KB 15 0
GRI 405: Diversity and Equal Opportunity 2016 gri (2016p) 2016 856 KB 15 1
GRI 406: Non-discrimination 2016 gri (2016q) 2016 853 KB 12 1
GRI 407: Freedom of Association and Collective Bargaining 2016 gri (2016r) 2016 870 KB 13 1
GRI 408: Child Labor 2016 gri (2016s) 2016 868 KB 14 0
GRI 409: Forced or Compulsory Labor 2016 gri (2016t) 2016 912 KB 13 1
GRI 410: Security Practices 2016 gri (2016u) 2016 841 KB 11 4
GRI 411: Rights of Indigenous Peoples 2016 gri (2016v) 2016 863 KB 14 1
GRI 413: Local Communities 2016 gri (2016w) 2016 885 KB 16 2
GRI 414: Supplier Social Assessment 2016 gri (2016x) 2016 853 KB 14 1
GRI 415: Public Policy 2016 gri (2016y) 2016 816 KB 12 1
GRI 416: Customer Health and Safety 2016 gri (2016z) 2016 825 KB 12 1
GRI 417: Marketing and Labeling 2016 gri (2016aa) 2016 837 KB 14 1
GRI 418: Customer Privacy 2016 gri (2016ab) 2016 825 KB 12 1
GRI Standards Glossary 2022 gri (2022ac) 2022 680 KB 23 2
A Practical Guide to Sustainability Reporting Using GRI and SASB Standards gri (2021ag) 2021 1.5 MB 42 4
The GRI Standards — A Guide for Policy Makers gri (2021af) 2021 7.1 MB 19 2
Total - 70.863MB 1,826 201
Table 4: Metadata for the GRI Standards collection in ESGenius. This comprehensive collection spans 2016-2024 and comprises: (1) Universal Standards (GRI 1-3) establishing core reporting principles, (2) Sector Standards (GRI 11-14) for high-impact industries, and (3) Topic Standards covering economic (200 series), environmental (300 series), and social (400 series) aspects. Key features include: Coverage: 1,826 pages across 40+ standards documents; Evaluation: 201 domain-specific questions; Scope: Comprehensive ESG disclosure requirements spanning corporate governance, environmental impact, and social responsibility. All standards sourced from GRI (https://www.globalreporting.org/), the leading authority in sustainability reporting frameworks.
Original Document Title Year Size Pages No. Qs
Apparel, Accessories & Footwear - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cg) 2023 410 KB 21 4
Appliance Manufacturing - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ch) 2023 340 KB 13 1
Building Products & Furnishings - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cn) 2023 369 KB 18 1
E-commerce - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cy) 2023 402 KB 24 3
Household & Personal Products - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dn) 2023 371 KB 19 1
Multiline & Specialty Retailers & Distributors - SUSTAINABILITY ACCOUNTING STANDARD sas (2023eb) 2023 399 KB 24 3
Toys & Sporting Goods - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ev) 2023 346 KB 13 0
Coal Operations - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cr) 2023 490 KB 41 8
Construction Materials - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ct) 2023 414 KB 27 3
Iron & Steel Producers - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ds) 2023 392 KB 24 4
Metals & Mining - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dz) 2023 521 KB 47 8
Oil & Gas – Exploration & Production - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ed) 2023 511 KB 47 6
Oil & Gas – Midstream - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ee) 2023 407 KB 26 4
Oil & Gas – Refining & Marketing - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ef) 2023 433 KB 31 5
Oil & Gas – Services - SUSTAINABILITY ACCOUNTING STANDARD sas (2023eg) 2023 413 KB 28 4
Asset Management & Custody Activities - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ci) 2023 404 KB 24 3
Commercial Banks - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cs) 2023 394 KB 23 3
Consumer Finance - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cu) 2023 372 KB 20 3
Insurance - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dp) 2023 408 KB 26 3
Investment Banking & Brokerage - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dr) 2023 420 KB 28 4
Mortgage Finance - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ea) 2023 366 KB 17 2
Security & Commodity Exchanges - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ep) 2023 362 KB 17 1
Agricultural Products - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cc) 2023 426 KB 30 3
Alcoholic Beverages - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cf) 2023 400 KB 26 4
Food Retailers & Distributors - SUSTAINABILITY ACCOUNTING STANDARD sas (2023de) 2023 452 KB 37 5
Meat, Poultry & Dairy - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dw) 2023 442 KB 34 6
Non-Alcoholic Beverages - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ec) 2023 419 KB 30 3
Processed Foods - SUSTAINABILITY ACCOUNTING STANDARD sas (2023eh) 2023 434 KB 32 5
Restaurants - SUSTAINABILITY ACCOUNTING STANDARD sas (2023en) 2023 418 KB 30 5
Tobacco - SUSTAINABILITY ACCOUNTING STANDARD sas (2023eu) 2023 337 KB 12 1
Biotechnology & Pharmaceuticals - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cm) 2023 415 KB 29 4
Drug Retailers - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cx) 2023 378 KB 21 3
Health Care Delivery - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dj) 2023 432 KB 33 6
Health Care Distributors - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dk) 2023 358 KB 17 1
Managed Care - SUSTAINABILITY ACCOUNTING STANDARD sas (2023du) 2023 371 KB 20 3
Medical Equipment & Supplies - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dy) 2023 380 KB 22 2
Electric Utilities & Power Generators - SUSTAINABILITY ACCOUNTING STANDARD sas (2023da) 2023 458 KB 35 4
Engineering & Construction Service - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dd) 2023 402 KB 26 4
Table 5: Metadata for SASB industry-specific disclosure standards — Part I.
Original Document Title Year Size Pages No. Qs
Gas Utilities & Distributors - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dh) 2023 372 KB 19 0
Home Builders - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dl) 2023 383 KB 21 1
Real Estate - SUSTAINABILITY ACCOUNTING STANDARD sas (2023el) 2023 285 KB 38 5
Real Estate Services - SUSTAINABILITY ACCOUNTING STANDARD sas (2023em) 2023 349 KB 15 0
Waste Management - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ew) 2023 430 KB 31 5
Water Utilities & Services - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ex) 2023 436 KB 32 4
Biofuels - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cl) 2023 380 KB 20 1
Forestry Management - SUSTAINABILITY ACCOUNTING STANDARD sas (2023df) 2023 363 KB 18 2
Fuel Cells & Industrial Batteries - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dg) 2023 381 KB 21 3
Pulp & Paper Products - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ej) 2023 390 KB 24 2
Solar Technology & Project Developers - SUSTAINABILITY ACCOUNTING STANDARD sas (2023es) 2023 410 KB 28 5
Wind Technology & Project Developers - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ey) 2023 358 KB 17 2
Aerospace & Defence - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cb) 2023 439 KB 27 4
Chemicals - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cq) 2023 447 KB 36 7
Containers & Packaging - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cv) 2023 422 KB 30 5
Electrical & Electronic Equipment - SUSTAINABILITY ACCOUNTING STANDARD sas (2023db) 2023 385 KB 23 3
Industrial Machinery & Goods - SUSTAINABILITY ACCOUNTING STANDARD sas (2023do) 2023 361 KB 17 2
Advertising & Marketing - SUSTAINABILITY ACCOUNTING STANDARD sas (2023) 2023 368 KB 18 1
Casinos & Gaming - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cp) 2023 353 KB 15 1
Education - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cz) 2023 366 KB 18 1
Hotels & Lodging - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dm) 2023 371 KB 19 2
Leisure Facilities - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dt) 2023 343 KB 13 1
Media & Entertainment - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dx) 2023 357 KB 15 1
Professional & Commercial Services - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ei) 2023 371 KB 18 2
Electronic Manufacturing Services & Original Design Manufacturing - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dc) 2023 379 KB 22 3
Hardware - SUSTAINABILITY ACCOUNTING STANDARD sas (2023di) 2023 392 KB 24 3
Internet Media & Services - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dq) 2023 417 KB 27 4
Semiconductors - SUSTAINABILITY ACCOUNTING STANDARD sas (2023eq) 2023 409 KB 27 4
Software & IT Services - SUSTAINABILITY ACCOUNTING STANDARD sas (2023er) 2023 426 KB 29 4
Telecommunication Services - SUSTAINABILITY ACCOUNTING STANDARD sas (2023et) 2023 409 KB 26 2
Air Freight & Logistics - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cd) 2023 394 KB 24 4
Airlines - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ce) 2023 371 KB 19 2
Auto Parts - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cj) 2023 373 KB 20 2
Automobiles - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ck) 2023 378 KB 22 3
Car Rental & Leasing - SUSTAINABILITY ACCOUNTING STANDARD sas (2023co) 2023 334 KB 11 3
Cruise Lines - SUSTAINABILITY ACCOUNTING STANDARD sas (2023cw) 2023 414 KB 26 3
Marine Transportation - SUSTAINABILITY ACCOUNTING STANDARD sas (2023dv) 2023 400 KB 24 2
Rail Transportation - SUSTAINABILITY ACCOUNTING STANDARD sas (2023ek) 2023 381 KB 21 3
Road Transportation - SUSTAINABILITY ACCOUNTING STANDARD sas (2023eo) 2023 368 KB 17 1
Total 30.431MB 1,864 236
Table 6: Comprehensive metadata for all 77 industry-specific SASB Sustainability Accounting Standards issued in 2023 and now stewarded by the IFRS Foundation’s ISSB. This corpus, distilled for our ESGenius system, comprises 77 documents totalling 30.431 MB, 1,864 pages, and 236 MCQs. The standards span the full economy—from Apparel to Transportation—and cover financially material sustainability topics including greenhouse gas emissions, water management, data security, and workforce diversity. These standards provide investors with decision-relevant ESG information for valuation, risk assessment, and stewardship. The complete collection is freely accessible at https://www.sasb.org/.
Original Document Title Year Size Pages No. Qs
IFRS S1 ifr (2023ad) 2023 307 KB 48 8
IFRS S2 ifr (2023af) 2023 297 KB 46 8
Progress on Corporate Climate-related Disclosures — 2024 Report ifr (2024cb) 2024 1.3 MB 164 27
IFRS S1 Basis for Conclusions on General Requirements for Disclosure of Sustainability-related Financial Information ifr (2023ac) 2023 341 KB 57 10
IFRS S2 Basis for Conclusions on Climate-related Disclosures ifr (2023ae) 2023 337 KB 55 11
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures ifr (2023ag) 2023 2.3 MB 538 33
IFRS S1 Accompanying Guidance on General Requirements for Disclosure of Sustainability-related Financial Information ifr (2023ab) 2023 201 KB 17 2
Exposure Draft Amendments to Greenhouse Gas Emissions Disclosures Proposed amendments to IFRS S2 Comments to be received by 27 June 2025 ifr (2025) 2025 239 KB 24 4
Exposure Draft Basis for Conclusions on Amendments to Greenhouse Gas Emissions Disclosures Proposed amendments to IFRS S2 Comments to be received by 27 June 2025 ifr (2025cb) 2025 238 KB 25 4
Comparison IFRS S2 Climate-related Disclosures with the TCFD recommendations ifr (2023) 2024 105 KB 12 1
IFRS Taxonomy IFRS Sustainability Disclosure Taxonomy 2024 IFRS S1 General Requirements for Disclosure of Sustainability-related Financial Information and IFRS S2 Climate-related Disclosures ifr (2024) 2024 1.1 MB 76 10
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 1 — Apparel, Accessories & Footwear ifr (2023ah) 2023 194 KB 10 4
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 2 — Appliance Manufacturing ifr (2023as) 2023 179 KB 7 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 3 — Building Products & Furnishings ifr (2023bd) 2023 203 KB 13 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 4 — E-Commerce ifr (2023bo) 2023 199 KB 11 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 5 — Household & Personal Products ifr (2023bz) 2023 191 KB 9 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 6 — Multiline and Specialty Retailers & Distributors ifr (2023ck) 2023 180 KB 7 0
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 7 — Coal Operations ifr (2023cu) 2023 216 KB 14 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 8 — Construction Materials ifr (2023cv) 2023 232 KB 17 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 9 — Iron & Steel Producers ifr (2023cw) 2023 210 KB 14 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 10 — Metals & Mining ifr (2023ai) 2023 204 KB 12 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 11 — Oil & Gas – Exploration & Production ifr (2023aj) 2023 244 KB 20 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 12 — Oil & Gas – Midstream ifr (2023ak) 2023 196 KB 10 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 13 — Oil & Gas – Refining & Marketing ifr (2023al) 2023 212 KB 13 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 14 — Oil & Gas – Services ifr (2023am) 2023 202 KB 11 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 15 — Asset Management & Custody Activities ifr (2023an) 2023 202 KB 11 4
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 16 — Commercial Banks ifr (2023ao) 2023 189 KB 8 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 17 — Insurance ifr (2023ap) 2023 223 KB 15 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 18 — Investment Banking & Brokerage ifr (2023aq) 2023 201 KB 10 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 19 — Mortgage Finance ifr (2023ar) 2023 188 KB 7 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 20 — Agricultural Products ifr (2023at) 2023 223 KB 17 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 21 — Alcoholic Beverages ifr (2023au) 2023 209 KB 13 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 22 — Food Retailers & Distributors ifr (2023av) 2023 219 KB 15 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 23 — Meat, Poultry & Dairy ifr (2023aw) 2023 235 KB 19 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 24 — Non-Alcoholic Beverages ifr (2023ax) 2023 217 KB 15 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 25 — Processed Foods ifr (2023ay) 2023 219 KB 15 0
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 26 - Restaurants ifr (2023az) 2023 202 KB 11 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 27 — Drug Retailers ifr (2023ba) 2023 186 KB 7 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 28 — Health Care Delivery ifr (2023bb) 2023 201 KB 11 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 29 — Health Care Distributors ifr (2023bc) 2023 180 KB 6 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 30 — Managed Care ifr (2023be) 2023 174 KB 6 1
Table 7: IFRS / ISSB sustainability disclosure materials — Part I.
Original Document Title Year Size Pages No. Qs
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 31 — Medical Equipment & Supplies ifr (2023bf) 2023 183 KB 7 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 32 — Electric Utilities & Power Generators ifr (2023bg) 2023 257 KB 24 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 33 — Engineering & Construction Services ifr (2023bh) 2023 222 KB 16 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 34 — Gas Utilities & Distributors ifr (2023bi) 2023 210 KB 13 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 35 — Home Builders ifr (2023bj) 2023 214 KB 13 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 36 — Real Estate ifr (2023bk) 2023 285 KB 34 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 37 — Real Estate Services ifr (2023bl) 2023 197 KB 9 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 38 — Waste Management ifr (2023bm) 2023 215 KB 12 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 39 — Water Utilities & Services ifr (2023bn) 2023 241 KB 19 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 40 — Biofuels ifr (2023bp) 2023 208 KB 14 0
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 41 — Forestry Management ifr (2023bq) 2023 209 KB 12 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 42 — Fuel Cells & Industrial Batteries ifr (2023br) 2023 202 KB 10 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 43 — Pulp & Paper Products ifr (2023bs) 2023 202 KB 18 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 44 — Solar Technology & Project Developers ifr (2023bt) 2023 216 KB 14 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 45 — Wind Technology & Project Developers ifr (2023bu) 2023 195 KB 8 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 46 — Aerospace & Defence ifr (2023bv) 2023 204 KB 9 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 47 — Chemicals ifr (2023bw) 2023 224 KB 16 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 48 — Containers & Packaging ifr (2023bx) 2023 232 KB 17 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 49 — Electrical & Electronic Equipment ifr (2023by) 2023 202 KB 10 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 50 — Industrial Machinery & Goods ifr (2023ca) 2023 196 KB 9 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 51 — Casinos & Gaming ifr (2023cb) 2023 189 KB 7 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 52 — Hotels & Lodging ifr (2023cc) 2023 201 KB 9 4
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 53 — Leisure Facilities ifr (2023cd) 2023 188 KB 7 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 54 — Electronic Mfg Services & ODM ifr (2023ce) 2023 196 KB 8 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 55 — Hardware ifr (2023cf) 2023 200 KB 10 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 56 — Internet Media & Services ifr (2023cg) 2023 200 KB 9 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 57 — Semiconductors ifr (2023ch) 2023 222 KB 14 0
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 58 — Software & IT Services ifr (2023ci) 2023 208 KB 11 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 59 — Telecommunication Services ifr (2023cj) 2023 205 KB 10 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 60 — Air Freight & Logistics ifr (2023cl) 2023 202 KB 11 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 61 — Airlines ifr (2023cm) 2023 201 KB 9 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 62 — Auto Parts ifr (2023cn) 2023 192 KB 8 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 63 — Automobiles ifr (2023co) 2023 193 KB 8 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 64 — Car Rental & Leasing ifr (2023cp) 2023 189 KB 7 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 65 — Cruise Lines ifr (2023cq) 2023 206 KB 10 3
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 66 — Marine Transportation ifr (2023cr) 2023 208 KB 10 2
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 67 — Rail Transportation ifr (2023cs) 2023 201 KB 9 1
IFRS S2 Industry-based Guidance on implementing Climate-related Disclosures Vol 68 — Road Transportation ifr (2023ct) 2023 200 KB 9 3
Total - 20.81MB 1,866 238
Table 8: Comprehensive metadata for the entire IFRS / ISSB sustainability-disclosure corpus (2023–2025): it comprises the universal core standards IFRS S1 (general sustainability-related financial disclosure) and IFRS S2 (climate-related disclosure), their bases for conclusions, accompanying guidance, the 2024 climate-progress report, the 2024 Sustainability Disclosure Taxonomy, the 2025 greenhouse-gas exposure draft (with basis), a comparison of IFRS S2 with the TCFD recommendations, and the 68-volume IFRS S2 Industry-based Guidance that maps SASB’s sector-specific materiality into the new framework. Together the 77 PDFs, total 20.81 MB, span 1,866 pages—well over two thousand when including ancillary matter—and yield 238 benchmark MCQs for ESGenius. IFRS S1 sets the universal disclosure baseline, while IFRS S2 details climate-specific metrics, mirroring the TCFD architecture and enriched with SASB’s sectoral depth; the accompanying materials establish a globally consistent, investor-focused baseline that links decision-relevant sustainability information—such as greenhouse-gas emissions, transition plans, climate resilience, data security, and workforce diversity—directly to financial statements. All documents are freely available at https://www.ifrs.org/.
Original Document Title Year Size Pages No. Qs
Task Force on Climate-related Financial Disclosures Overview tcf (2022) 2022 11.1 MB 25 2
Task Force on Climate-related Financial Disclosures Guidance on Metrics, Targets, and Transition Plans tcf (2021) 2021 12 MB 79 11
Task Force on Climate-related Financial Disclosures Implementing the Recommendations of the Task Force on Climate-related Financial Disclosures tcf (2021eb) 2021 1.1 MB 88 18
TCFD Workshop – Session 1: Fundamentals and Overview of TCFD tcf (2022eb) 2022 3.2 MB 40 7
TCFD Workshop – Session 2: Governance tcf (2022ec) 2022 1.4 MB 20 1
TCFD Workshop – Session 3: Strategy tcf (2022ed) 2022 2.7 MB 43 8
TCFD Workshop – Session 4: Risk Management tcf (2022ee) 2022 1.4 MB 34 4
TCFD Workshop – Session 5: Metrics and Targets tcf (2022ef) 2022 3 MB 53 9
Task Force on Climate-related Financial Disclosures Guidance on Scenario Analysis for Non-Financial Companies tcf (2020eb) 2020 3.7 MB 133 22
Task Force on Climate-related Financial Disclosures Guidance on Risk Management Integration and Disclosure tcf (2020) 2020 5.2 MB 52 9
Task Force on Climate-related Financial Disclosures 2023 Status Report tcf (2023) 2023 19.3 MB 161 32
Total - 60 MB 728 123
Table 9: Comprehensive metadata for the Task Force on Climate-related Financial Disclosures (TCFD) corpus included in ESGenius. The collection spans 2020–2023 and comprises 11 documents totaling 60 MB across 728 pages, yielding 123 benchmark questions. The materials include core guidance documents on implementation, metrics, targets, scenario analysis and risk management, a complete five-part workshop series covering fundamentals through metrics, and the 2023 status report. These documents establish the foundation for global climate-risk disclosure practices and form the conceptual framework adopted by ISSB’s IFRS S2. All materials are accessible at https://www.fsb-tcfd.org/.
Original Document Title Year Size Pages No. Qs
CDP Full Corporate Questionnaire April 2025 – Modules 1–6 cdp (2025c) 2025 5.1 MB 447 29
CDP Full Corporate Questionnaire April 2025 – Module 7 cdp (2025b) 2025 4.9 MB 482 41
CDP Full Corporate Questionnaire April 2025 – Modules 8–13 cdp (2025d) 2025 4.7 MB 435 44
CDP SME Questionnaire April 2025 – Modules 14–21 cdp (2025e) 2025 2.1 MB 174 38
2025 CDP–ICLEI Track and States & Regions Questionnaire and Guidance cdp (2025a) 2025 2.5 MB 430 31
CDP Full Corporate Scoring Introduction 2024 cdp (2024) 2024 293 KB 19 2
Total - 19.6 MB 1,987 185
Table 10: Comprehensive metadata for the Carbon Disclosure Project (CDP) knowledge base curated in ESGenius: the collection comprises six hyper-linked PDFs including the 2025 Full Corporate Questionnaire split across three modules (114 questions, 1,364 pages), the 2025 SME Questionnaire (38 questions, 174 pages), the 2025 CDP–ICLEI Track and States & Regions Questionnaire (31 questions, 430 pages), and the 2024 Full Corporate Scoring Introduction (2 scoring criteria, 19 pages). Together the corpus totals 19.6 MB and spans 1,987 pages, delivering 185 unique, standardised questions that power the world’s largest voluntary platform for climate, water-security, and deforestation disclosure. All files are directly accessible via the linked titles, with consolidated resources available at https://www.cdp.net/.
Original Document Title Year Size Pages No. Qs
The 2030 Agenda for Sustainable Development’s 17 Sustainable Development Goals (SDGs) sdg (2015) 2015 424 KB 19 1
Transforming Our World: The 2030 Agenda for Sustainable Development tra (2015) 2015 378 KB 41 4
Progress towards the Sustainable Development Goals — Report of the Secretary-General sdg (2024) 2024 518 KB 26 2
The Sustainable Development Goals Report 2024 sdg (2024eb) 2024 8.6 MB 51 6
Sustainable Development Goals – Briefing Book 2023 (UN Office for Partnerships) sdg (2023) 2023 4.6 MB 35 7
GAR Special Report 2023 — Mapping Resilience for the Sustainable Development Goals gar (2023) 2023 19.5 MB 51 5
Sustainable Development Report 2024 — The SDGs and the UN Summit of the Future (includes the SDG Index and Dashboards) sdr (2024) 2024 39.3 MB 512 12
Total - 73.33 MB 735 37
Table 11: Comprehensive metadata for the United Nations Sustainable Development Goals (SDGs) corpus incorporated in ESGenius: the set begins with the 2015 adoption texts—the 19-page plain-language overview of the 17 Goals and the 41-page General Assembly resolution “Transforming Our World” that enshrines the 2030 Agenda—then follows implementation through the Secretary-General’s Progress towards the SDGs 2024 and the flagship SDGs Report 2024, is complemented by the UN Office for Partnerships’ SDG Briefing Book 2023 and deepened by thematic analyses such as UNDRR’s 2023 Global Assessment Report on Resilience (GAR) and the independent Sustainable Development Report 2024 with its widely cited SDG Index and global dashboards. Across these seven key PDFs—totalling 73.33 MB, 735 pages, and distilled into 37 benchmark MCQs that anchors national, corporate, and investor sustainability strategies in the universally agreed 17-goal framework. Full SDG resources are freely available at https://sdgs.un.org/goals.
Figure 4: This cloud visualizes the 1,136 question stems after filtering generic fillers. Dominant terms such as “disclosures”, “climate-related”, “sustainability”, “accurately”, and “IFRS” reveal that the questions emphasize reporting frameworks and precision in interpreting ESG guidance. The prominence of verbs like “described”, “outlined”, and “reflects” indicates a consistent demand for higher-order reasoning (e.g., identifying relationships, implications, or best interpretations rather than simple fact recall).
Figure 5: The aggregate vocabulary of the 5,680 answer choices centers on the same core ESG-reporting nouns that dominate the source text—“emissions,” “entity,” “organization,” “energy,” “water,” “risks,” and “reporting”—but it is studded with decisive qualifiers such as “without,” “due,” “exclusively,” “primarily,” and “regardless.” These modifiers reveal how distractors are engineered: they adjust scope, responsibility, or conditionality to make each option plausible while still allowing only one to satisfy the nuanced criteria posed by the question.
Figure 6: Drawn from the reference excerpts linked to each question, this word cloud highlights the knowledge backbone behind the benchmark. Key nouns—“entity”, “organization”, “emissions”, “energy”, “water”, “reporting”, “standards”—underscore the dataset’s strong focus on corporate disclosure obligations, accounting boundaries (Scope 1/2/3 GHG), and resource-specific metrics. Frequent technical modifiers such as “financial”, “material”, “percentage”, “management” show that passages often quantify impacts or prescribe measurement criteria, aligning with the analytical depth expected of the MCQs.
Figure 7: Relative question distribution (benchmark MCQs) derived from each source family. In contrast to the page distribution, questions are more evenly spread: the largest contributors are IFRS/ISSB (20.9%, 237 Qs) and SASB (20.8%, 235 Qs), followed by GRI (17.7%, 201 Qs) and CDP (16.3%, 184 Qs). TCFD (10.8%, 122 Qs) and IPCC (10.3%, 116 Qs) provide focused climate-risk and climate-science coverage, while the SDG set supplies a compact but essential sustainability anchor (3.3%, 36 Qs).
Figure 8: Relative page‐count distribution of the ESGenius-Corpus across its 7 source families. The Intergovernmental Panel on Climate Change (IPCC) alone accounts for a majority of the material—54%, or 10,582 pages—reflecting the encyclopaedic scope of its assessment reports. Standards and disclosure frameworks contribute smaller but still substantive shares: CDP (10.1%, 1,987 pp.), IFRS/ISSB (9.5%, 1,866 pp.), SASB (9.5%, 1,864 pp.), and GRI (9.3%, 1,826 pp.). Guidance‐oriented sources such as TCFD (3.7%, 728 pp.) and the UN SDGs corpus (3.8%, 735 pp.) round out the collection.
Figure 9: Example zero-shot performance heatmap showing model accuracy patterns across different ESG question types and topics within ESGenius (based on initial data). Darker colors indicate higher accuracy.
Figure 10: Conceptual information display for interactive heatmap cells (Figure 9), showing model and accuracy score.